solve the triangle in which c=3.4 cm , angleA=25,and angleB=85

Grade:12th pass

1 Answers

Raghu Vamshi Hemadri
8 years ago
From Sine rule
(SinA)/a=(SinB)/b=(SinC)/c
(Sin250)/a=(Sin850)/b=(Sin700)/3.4
a=1.5295
b=3.6054
c=3.4